A furniture manufacturer buys special new manufacturing equipment for 1.8 M$ that has an estimated salvage value of 0.3 M$ (M is mega-). If the useful life for tax purposes is 7 years, find the book value at the end of 4 years for MACRS depreciation.
Answer #1

Book value 4 year=$ 942,857.14

Explanation:

Depreciation=(cost -salvage)/useful life

=(1,800,000-30,000)/7

=1,500,000/7

=214,287.71

Depreciation for the 4 year=214,287.71*4

=857,142.86

Net book value=cost -accumulated depreciation

=1,800,000-857,142.86

=942,857.14